OpenStudy (anonymous):

solve by using tables. give each answer to at most two decimal places. -2x^2-4=-8 A. 0.59,3.41 B. -3.41,-0.59 C. 1.17, 6.83 D. -1.41,1.41

OpenStudy (azureilai):

You can just use quadratic formula for this

OpenStudy (anonymous):

but i dont understand how to do it :P

OpenStudy (anonymous):

check your question again

OpenStudy (anonymous):

i'm really not getting this :P

OpenStudy (anonymous):

wait a sec

OpenStudy (anonymous):

okie here we go.. we have equation -2x^2-4=-8 now add +8 to both side what we get -2x^2-4+8=-8+8 which is equal to -2x^2+4=0 now take -2 as a common factor from above so we get -2(x^2-2)=0 this gives x^2-2=0 which is in turn x^2=2 now taking square root of both sides we get x=+or-sqrt(2) from tables you can find square root of 2 so ans is -1.41 or 1.41

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

you are most welcome..